# A Modular-based Strategy for Mitigating Gradient Conflicts in Simultaneous Speech Translation

### v2

@@ Abstract
One of the most common approaches to enhance SimulST model performance is Multi-task Learning (MTL)[7], which integrates auxiliary tasks such as Simultaneous Automatic Speech Recognition (SimulASR)[8] and Simultaneous Machine Translation (SimulMT)[9]. These auxiliary tasks provide additional contextual information, which helps improve the performance of the primary task. While MTL has demonstrated advantages in enhancing overall model effectiveness, it also introduces substantial challenges, particularly the potential for the conflicts between task optimizations.
@@ II Method
In this section, we formalize the SimulST task within the MTL framework and provide a detailed explanation of the method. We begin by analyzing the limitations of resolving gradient conflicts at the model level, followed by the introduction of our proposed modular gradient conflict mitigation strategy.
II-A SimulST in MTL
Giving the training data as a triple , denotes the acoustic features extracted from the input speech in the source language, denotes the corresponding transcription, and denotes the text in the target language. The learning objective for training is as follows:
where , , and are the cross-entropy loss function of pairs s y, s x and x y.
II-B Limitations of Model-level Gradient Conflict Resolution
@@
II-B1 Gradient Conflict Masking
We monitor a contradicting situation in that PCGrad often computes near-zero cosine similarity, suggesting that task optimization directions are nearly orthogonal [12, 13, 14]. Further analysis shows that gradient conflicts are usually confined to specific model components like attention mechanisms, and become evident only after extended optimization. Unfortunately, these conflicts are usually masked because when calculating cosine similarity via such a large model-level gradient vector, more contributions from the non-conflicting modules are dominating.
II-B2 GPU Memory Consumption
@@
II-C MGCM: Modular Gradient Conflict Mitigation
@@ II-C1 Modularization
II-C2 Conflict Detection and Projection
@@
The model parameters are then updated utilizing these adjusted gradients, minimizing conflicts at the module level. By focusing on module-level conflicts, MGCM effectively addresses gradient conflict masking observed in PCGrad. Meanwhile, the modular approach leads to an incidental benefit of GPU memory saving, as it avoids the need to compute long vectors. The resulting memory savings are quantified and discussed in Section III-F.
@@ III-E Conflict Probability Analysis
III-F GPU Memory Consumption

### v3

@@ Abstract
One of the most common approaches to enhance SimulST model performance is Multi-task Learning (MTL)[7], which integrates auxiliary tasks such as Simultaneous Automatic Speech Recognition (SimulASR)[8, 9] and Simultaneous Machine Translation (SimulMT)[10]. These auxiliary tasks provide additional contextual information, which helps improve the performance of the primary task. While MTL has demonstrated advantages in enhancing overall model effectiveness, it also introduces substantial challenges, particularly the potential for the conflicts between task optimizations.
@@ II Method
In this section, we provide a detailed explanation of the method. We begin by analyzing the limitations of resolving gradient conflicts at the model level, followed by the introduction of our proposed modular gradient conflict mitigation strategy.
II-A Limitations of Model-level Gradient Conflict Resolution
@@
II-A1 Gradient Conflict Masking
We monitor a contradicting situation in that PCGrad often computes near-zero cosine similarity, suggesting that task optimization directions are nearly orthogonal [13, 14, 15]. Further analysis shows that gradient conflicts are usually confined to specific model components like attention mechanisms, and become evident only after further optimization. Unfortunately, these conflicts are usually masked because when calculating cosine similarity via such a large model-level gradient vector, more contributions from the non-conflicting modules are dominating.
II-A2 GPU Memory Consumption
@@
II-B MGCM: Modular Gradient Conflict Mitigation
@@ II-B1 Modularization
II-B2 Conflict Detection and Projection
@@
The model parameters are then updated utilizing these adjusted gradients, minimizing conflicts at the module level. By focusing on module-level conflicts, MGCM effectively addresses gradient conflict masking observed in PCGrad. Meanwhile, the modular approach leads to an incidental benefit of GPU memory saving, as it avoids the need to compute long vectors. The resulting memory savings are quantified and discussed in Section III-G.
@@ III-E Conflict Probability Analysis
III-F Statistical Significance Test
We further ensure the reliability of MGCM by employing statistical significance tests. When the p-value is greater than 0.05, it indicates that the observed differences between the methods are probably due to randomness. We perform tests between online and offline inference settings and find that the maximum p-value is 0.034 in En-De. All values are well below the 0.05 threshold, ensuring that the performance improvements achieved by MGCM are statistically significant and not due to random fluctuations.
III-G GPU Memory Consumption